# In-Context Freeze-Thaw Bayesian Optimization for Hyperparameter Optimization

### v1

@@ A.2 Meta-training Data Generating Procedure
A single meta-training example in our setting corresponds to a training set $D_{train}$ and test set $D_{test}$, where $D_{train}=\{(\lambda,t),f(\lambda,t)\}_{t=1,\forall\lambda\in\Lambda}^{b_{\lambda}^{\prime}}$} corresponds to the (synthetic) partial learning curves observed thus far (i.e., the analog of $H$ at test time) and $D_{test}\subseteq{\lambda,t,\,f(\lambda,t)\}_{\lambda\in\Lambda,\,b_{max}\geq t>b_{\lambda}}}$ the extrapolation targets we want FT-PFN to predict. To keep the input size of FT-PFN fixed we choose $|D_{train}|+|D_{test}|=N$ and vary the size of $|D_{train}|\sim\mathcal{U}(0,B-1)$. As $b_{max}$ varies in practice, we sample it log-uniformly in $[1,1000]$. Note that in the special case $b_{max}=1$, we train FT-PFN for $B$-BO. $\Lambda={\lambda_{i}}_{i}=1^{N}$ is our synthetic configuration space with $\lambda_{i}\sim\mathcal{U}(0,1)^{d}$, with $|\lambda_{i}|=m\sim\mathcal{U}(0,M)$ the dimensionality of our configuration space. We determine $b_{\lambda}$ by sampling a bag of $|D_{train}|$ elements from $\Lambda$ proportionally to weights $\{w_{\lambda}\}_{\lambda\in\Lambda}$ that follow a Dirichlet distribution with $\log_{10}(\alpha)\sim\mathcal{U}(-4,-1)$ resulting in heterogeneous budget allocations that vary from breadth-first to depth-first. [^3] We adopt the same strategy to generate benchmark tasks for our evaluation of prediction quality described in Section 5.1. We use the same weights to sample another bag of $|D_{test}|$ determining the number of extrapolation targets for each $\lambda$, where each target is chosen $\mathcal{U}(b_{\lambda},b_{max})$. Finally, to generate the corresponding $f(\lambda_{i},t_{i})$, we first instantiate the random variables that are task-specific but do not depend on $\lambda$, i.e., $y_{0}\sim\mathrm{Beta}(1,2)$, and the architecture and weights $\theta$ of the neural network pp. Finally, we determine $f(\lambda_{i},t_{i})=\hat{f}(t_{i},p(\lambda;\theta))$.
@@ A.3 Architecture and Hyperparameters
Following Müller et al. (2022), we use a sequence Transformer (Vaswani et al., 2017) for FT-PFN and treat each tuple ($\lambda$, $t$, f($\lambda$, $t$)) (for train) and ($\lambda$, $t$) (for test) as a separate position/token. We do not use positional encoding such that we are permutation invariant. FT-PFN outputs a discretized approximation of the PPD, each output corresponding to the probability density of one of the equal-sized bins. We set the number of bins/outputs to 1,000. For the transformer, we use 6 layers, an embedding size of 512, four heads, and a hidden size of 1,024, resulting in a total of 14.69M parameters. We use a standard training procedure for all experiments, minimizing the cross-entropy loss from Equation 1 on 2.0M synthethic datasets generated as described in Section A.2, using the Adam optimizer (Kingma et al., 2015) (learning rate 0.0001, batch size 25) with cosine annealing (Loshchilov & Hutter, 2017) [Loshchilov and Hutter, 2017] with a linear warmup over the first 25% epochs of the training. Training took roughly 8 GPU hours on an RTX2080 GPU and the same FT-PFN is used in all experiments described in Section 5, without any retraining/fine-tuning.
@@ A.4 Acquisition function
Algorithm 2 describes the acquisition loop in ifBO. The two parameters in Equation 4 are chosen randomly, with Probability Improvement (PI) (Mockus et al., 1978) as the base acquisition function (AF). In each iteration of ifBO (L3-L9 in Alg. 1), Alg. 2 is invoked once. In each such run, the random horizon $z$ and the scaled threshold of improvement $f_{\text{threshold}}$ are sampled once. This process can be seen as instantiating an acquisition function from a portfolio of multi-fidelity PIs. The choice of PI, the multi-fidelity component of extrapolating hyperparameters, and the random selection of an acquisition behaviour lends the naming of this acquisition function, MFPI-random. For each candidate evaluated ($\hat{\lambda}\in\Lambda$), the posterior is inferred given the surrogate for a hyperparameter $\hat{\lambda}$ at a total horizon of its current observed length in unit steps ($b_{\hat{\lambda}}$) ++ the sampled horizon ($z$). The candidate with the highest obtained PI score is returned as the candidate solution to query next in the main Alg. 1 loop.
@@
maximal steps $b_{\text{max}}$
Output: $\lambda^{\prime}\in\Lambda$, next query
@@
  $f_{\text{best}}\leftarrow$ best score seen in $H$
  $z\sim\mathcal{U}(1,b_{\text{max}})$random horizon
  $\tau\sim\mathcal{U}(-4,-1)$random threshold scaling
  $f_{\text{threshold}}=f_{\text{best}}\cdot(1-10^{\tau})$
  $D\leftarrow\{\}$
  for all $\hat{\lambda}\in\Lambda$ do
   $y^{\prime}\leftarrow$ $\mathds{P}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}(\hat{\lambda},~b_{\hat{\lambda}}+z)\leq f_{\text{threshold}})$
   $D\leftarrow{}D\bigcup{}\{(\hat{\lambda},~y^{\prime})\}$
  end for
  return $\operatorname*{arg\!max}_{\lambda\in\Lambda}\{(\lambda,y)\}_{(\lambda,y)\in D}$
@@ Appendix B Benchmarks
This choice of benchmarks is largely dictated by following the existing benchmarks used in prior work, especially the two primary baselines with which we compare to, DyHPO and DPL.
These benchmarks were provided using mf-prior-bench [^4] https://github.com/automl/mf-prior-bench.
@@
    Each benchmark ranges in the size of their learning curves, depending on the task, ranging from 5 to $14$. For each task, there are different variant based on a pair of dataset and batchsize. In total we evaluate our method on the 16 PD1 tasks below.
@@
Appendix C Further Ablations
C.1 Effectiveness of modeling curve divergence
@@
C.2 Pairwise comparison of freeze-thaw approaches
@@
C.3 Acquisition function ablation of the baselines
In this section, our objective is to explore the impact of incorporating randomization into the acquisition function on the baseline methods (DPL and DyHPO). For this purpose, we assess each baseline across four distinct acquisition functions (Figure 8), each representing a different configuration of Equation 4. The variants include: (ours), where both the horizon and the threshold for improvement are randomly selected, similar to the approach in ifBO; (one-step), where the horizon and threshold for improvement are chosen as in DyHPO; (at max), where the selection criteria for the horizon and threshold follow the methodology in DPL; and (random horizon), where the horizon is randomly determined, and the threshold is set to the best value observed.
@@ C.3 Acquisition function ablation of the baselines
Figure 8: Relative ablation over the horizon and threshold parameters of a multi-fidelity AF (Equation 4).
C.4 Comparison of freeze-thaw approaches with MFPI-random
In Figure 9, we present a comparison of all the freeze-thaw approaches—including ours, DPL, and DyHPO—when employing our acquisition function (MFPI-random). Despite all models utilizing the same acquisition function, our model significantly outperforms those of DPL and DyHPO. This clearly indicates the crucial role our prior in achieving the final performance.
@@ C.4 Comparison of freeze-thaw approaches with MFPI-random
Appendix D Aggregate plots over time
@@
LCBench PD1 TaskSet Aggregated
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Comparing relative rank over wallclock time (in $s$) over different benchmark families and the aggregated result overall. ifBO is on average better than the baselines, DyHPO and DPL, except for the TaskSet benchmark family where DPL starts the best but ifBO improves with more budget. Plots have been cutoff at a common $10^{5}$ seconds for fair comparison.
Appendix E Per-task HPO Plots
In Section 3.1, we presented HPO results on each of these three benchmarks in a comprehensive form, averaging rank and normalized regrets across every task in the suite. These averages may hide / be susceptible to outliers. For completeness, Figures 11-16 provide regret plots for every task in the benchmark, averaged across the 10 seeds. We find that our method consistently performs on par, or better than the best previous best HPO method, especially in later stages of the search, without notable outliers.

### v2

@@ A.2 Meta-training Data Generating Procedure
A single meta-training example in our setting corresponds to a training set $D_{\text{train}}$ and test set $D_{\text{test}}$, where $D_{\text{train}}=\bigcup_{\lambda\in\Lambda}\bigl\{\bigl((\lambda,\frac{b}{\operatorname*{b_{\text{max}}}}),\operatorname*{\pi_{\text{curve}}}(\lambda,\frac{b}{\operatorname*{b_{\text{max}}}})\bigl)\bigl\}_{b=1}^{b_{\lambda}}$} corresponds to the (synthetic) partial learning curves observed thus far (i.e., the analog of $H$ at test time) and $D_{\text{test}}\subseteq\bigcup_{\lambda\in\Lambda}\{((\lambda,\frac{b}{\operatorname*{b_{\text{max}}}}),\operatorname*{\pi_{\text{curve}}}(\lambda,\frac{b}{\operatorname*{b_{\text{max}}}}))\}_{b=b_{\lambda}}^{\operatorname*{b_{\text{max}}}}$} the extrapolation targets we want FT-PFN to predict. To keep the input size of FT-PFN fixed we choose $|D_{\text{train}}|+|D_{\text{test}}|=N=1,000$ and vary the size of $|D_{\text{train}}|\sim\mathcal{U}(0,N-1)$. As $\operatorname*{b_{\text{max}}}$ varies in practice, we sample it log-uniformly in $[1,N]$. Note that in the special case $b_{\text{max}}=1$, we train FT-PFN for black box BO. $\Lambda=\{\lambda_{i}\}_{i=1}^{N}$ is our synthetic configuration space with $\lambda_{i}\sim\mathcal{U}(0,1)^{m}$, with $|\lambda_{i}|=m\sim\mathcal{U}(0,M)$ the dimensionality of our configuration space. We determine $b_{\lambda}$ by sampling a bag of $|D_{\text{train}}|$ elements from $\Lambda$ proportionally to weights $\{w_{\lambda}\}_{\lambda\in\Lambda}$ that follow a Dirichlet distribution with $\log_{10}(\alpha)\sim\mathcal{U}(-4,-1)$ resulting in heterogeneous budget allocations that vary from breadth-first to depth-first. [^3] We adopt the same strategy to generate benchmark tasks for our evaluation of prediction quality described in Section 5.1. We use the same weights to sample another bag of $|D_{\text{test}}|$ determining the number of extrapolation targets for each $\lambda$, where each target $b$ is chosen $\mathcal{U}(b_{\lambda},b_{\text{max}})$. Finally, to generate the corresponding performance observation/target, we first instantiate the random variables that are task-specific but do not depend on $\lambda$, i.e., $y_{0}$, $y_{\text{max}}$ and the architecture and weights $\theta$ of the neural network $\operatorname*{\pi_{\text{config}}}$; and subsequently obtain $\operatorname*{\pi_{\text{curve}}}(\lambda,\frac{b}{\operatorname*{b_{\text{max}}}})$ using Equation 2.
@@ A.3 Architecture and Hyperparameters
Following Müller et al. (2022), we use a sequence Transformer (Vaswani et al., 2017) for FT-PFN and treat each tuple ($\lambda$, $t$, $\operatorname*{\pi_{\text{curve}}}$($\lambda$, $t$)) (for train) and ($\lambda$, $t$) (for test) as a separate position/token. We do not use positional encoding such that we are permutation invariant. FT-PFN outputs a discretized approximation of the PPD, each output corresponding to the probability density of one of the equal-sized bins. We set the number of bins/outputs to 1,000. For the transformer, we use 6 layers, an embedding size of 512, four heads, and a hidden size of 1,024, resulting in a total of 14.69M parameters. We use a standard training procedure for all experiments, minimizing the cross-entropy loss from Equation 1 on 2.0M synthethic datasets generated as described in Section A.2, using the Adam optimizer (Kingma et al., 2015) (learning rate 0.0001, batch size 25) with cosine annealing (Loshchilov & Hutter, 2017) with a linear warmup over the first 25% epochs of the training. Training took roughly 8 GPU hours on an RTX2080 GPU and the same FT-PFN is used in all experiments described in Section 5, without any retraining/fine-tuning.
@@ A.4 Acquisition function
Algorithm 2 describes the acquisition procedure MFPI-random, used in ifBO. In each iteration of ifBO (L5-L11 in Algorithm 1), Algorithm 2 is invoked once taking as input the configuration space $\Lambda$, the surrogate model $\mathcal{M}$, the observed history $H$, and the maximal training steps $\operatorname*{b_{\text{max}}}$ of a configuration. First, the random horizon $h^{\text{rand}}$ and the scaled factor of improvement $\tau^{\text{rand}}$ (and thereby $T^{\text{rand}}$) are sampled once in every execution of the algorithm (L2-L3). This process can be seen as instantiating an acquisition function from a portfolio of multi-fidelity PIs. The choice of PI, the multi-fidelity component of extrapolating hyperparameters, and the random selection of an acquisition behaviour lends the naming of this acquisition function, MFPI-random. Then, for each candidate hyperparameter $\lambda\in\Lambda$, the performance of the hyperparameter at a total step of $b_{\lambda}+h^{\text{rand}}$ is inferred, using the surrogate $\mathcal{M}{}$. Finally, the candidate with the highest obtained PI score is returned as the candidate solution to query next in the main Algorithm 1 loop. Figure 6 illustrates the behavior of MFPI-random w.r.t some values of $h^{\text{rand}}$ and $T^{\text{rand}}$, with FT-PFN as a surrogate.
@@
   maximal steps $\operatorname*{b_{\text{max}}}$
Output: $\lambda\in\Lambda$, hyperparameter to evaluate next
@@
1:  $f_{\text{best}}\leftarrow\max\,\{y\}_{(\cdot,\cdot,y)\in H}$ best score seen in $H$
2:  $h^{\text{rand}}\sim\mathcal{U}(1,\operatorname*{b_{\text{max}}})$random horizon
3:  $T^{\text{rand}}=f_{\text{best}}+10^{\tau^{\text{rand}}}\cdot(1-f_{\text{best}})\quad\text{with}\quad\tau^{\text{rand}}\sim\mathcal{U}(-4,-1)$random threshold scaling
4:  return $\underset{\lambda\in\Lambda}{\operatorname*{arg\!max}}\,\mathds{P}(\mathcal{M}{}(\lambda,~\min(b_{\lambda}+h^{\text{rand}},\operatorname*{b_{\text{max}}}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1};H})>T^{\text{rand}})$ to perform in-context learning we pass $H$ as input to FT-PFN
Figure 6: Illustration of the MFPI acquisition (Equation 3). (Left) The figure shows a collection of partial learning curves and their corresponding continuations predicted by our FT-PFN model. Here again, we consider 3 hyperparameters whose values are mapped onto the color of the curves. (Right)The figure shows the color of the curve continued (i.e., maximizing MFPI) for different values of the horizon and threshold parameters. Note that the ranges shown (and scale used), match those sampled uniformly by MFPI-random (Equation 4) and consequently, the likelihood of continuing a specific curve is proportional to the surface area covered in this image by its corresponding color. Finally, note that the bright red color corresponds to starting a new curve.
@@ Appendix B Benchmarks
This choice of benchmarks is largely dictated by following the existing benchmarks used in prior work, especially the two primary baselines with which we compare to, DyHPO and DPL. These benchmarks were provided using mf-prior-bench [^4] https://github.com/automl/mf-prior-bench.
@@
    Each benchmark ranges in the size of their learning curves, depending on the task, ranging from 5 to 1414. For each task, there are different variant based on a pair of dataset and batchsize. In total we evaluate our method on the 16 PD1 tasks below.
@@
Appendix C Baselines
To use ifBO in practice for an HPO task, please refer to NePS [^5] https://automl.github.io/neps/latest/. All our baselines were developed into the NePS framework that we forked and copied into our setup. Below, we describe the basic configuration of these baselines that were included in our experiments.
All baseline implementations can be found under neps in our experiment code available at: https://github.com/automl/ifBO/tree/icml-2024.
C.1 General baselines
We chose random search based algorithms as baselines for the different benchmarks. This additionally also shows the utility of the different fidelity scheduling algorithms in HyperBand and ASHA which traverses the fidelity space in progressive geometric intervals, relying on strong performance correlation at these fidelity checkpoints. For these baselines, we chose the existing implementations in NePS, benchmarked in previously published work (Mallik et al., 2023b).
Random Search
Simply searches uniformly random in the hyperparameter space. The fidelity is set to the $b_{\text{max}}$ as specified by each benchmark instance (see, Appendix B). Therefore, as an example, a budget of $1000$ freeze-thaw steps, will be equivalent to $20$ full random search evaluations for LCBench and Taskset tasks.
HyperBand
The NePS implementation follows the algorithm described in Li et al. (2018) and uses the early stopping hyper-hyperparameter, as $\eta=3$. The $b_{\text{min}}$ is either $1$ or as specified by the benchmark instances. Similarly for $b_{\text{max}}$.
ASHA
The NePS implementation follows the algorithm described in Li et al. (2020b) and uses the early stopping hyper-hyperparameter, as $\eta=3$. The $b_{\text{min}}$ is either $1$ or as specified by the benchmark instances. Similarly for $b_{\text{max}}$.
C.2 Freeze-thaw baselines
Here we describe the set of freeze-thaw BO algorithms. We note that due to experimental framework (optimizer-benchmark interfacing and analysis) related differences, performing ablation studies on the original implementations of DyHPO and DPL were not straightforward. For consistency and reducing confounding factors, all experiments were performed with implementations in the same experimental framework. Each of the algorithms were implemented in our custom NePS framework.
Freeze-Thaw with GPs
This algorithm is designed to take one unit step per configuration in the fidelity space. The first $3$ samples are selected uniformly random, as influenced by the seed. Subsequently, a Gaussian Process (GP) is fit on the joint hyperparameter and fidelity space to predict the loss, as a surrogate model. This baseline uses the greedy MF-EI acquisition function from Wistuba et al. (2022). The GP here uses a standard 5/2-Matérn kernel with a lengthscale of $1.0$.
DyHPO
This implementation follows the exact details given in Wistuba et al. (2022) and their publicly available code [^6] https://github.com/releaunifreiburg/DyHPO/tree/main. For a quick hyper-hyperparameter glance, refer here: https://github.com/automl/ifBO/blob/icml-2024/src/pfns_hpo/pfns_hpo/configs/algorithm/dyhpo-neps-v2.yaml
DPL
This implementation follows the exact details given in Kadra et al. (2023) and their publicly available code [^7] https://github.com/releaunifreiburg/DPL/tree/main. For a quick hyper-hyperparameter glance, refer here: https://github.com/automl/ifBO/blob/icml-2024/src/pfns_hpo/pfns_hpo/configs/algorithm/dpl-neps-max.yaml
Appendix D Further Ablations
D.1 Effectiveness of modeling curve divergence
@@
D.2 Pairwise comparison of freeze-thaw approaches
@@
D.3 Acquisition function ablation of the baselines
In this section, our objective is to explore the impact of incorporating randomization into the acquisition function on the baseline methods (DPL and DyHPO). For this purpose, we assess each baseline across four distinct acquisition functions (Figure 9). The variants include: (ours), where both the horizon and the threshold for improvement are randomly selected, similar to the approach in ifBO; (one-step), where the horizon and threshold for improvement are chosen as in DyHPO; (at max), where the selection criteria for the horizon and threshold follow the methodology in DPL; and (random horizon), where the horizon is randomly determined, and the threshold is set to the best value observed.
@@ D.3 Acquisition function ablation of the baselines
Figure 9: Relative ablation over the horizon and threshold parameters of a multi-fidelity AF. For each algorithm, we take the AF designed into the original algorithm and ablate over the two variables: extrapolation horizon and the best performance threshold.
D.4 Comparison of freeze-thaw approaches with MFPI-random
In Figure 10, we present a comparison of freeze-thaw approaches—including ours, DPL, and DyHPO—when employing our acquisition function (MFPI-random). Despite all models utilizing the same acquisition function, our model significantly outperforms those of DPL and DyHPO. This clearly indicates the crucial role our prior in achieving the final performance.
@@ D.4 Comparison of freeze-thaw approaches with MFPI-random
Appendix E Aggregate plots over time
@@
Figure 11: Comparing relative rank over wallclock time (in $s$) over different benchmark families and the aggregated result overall. ifBO is on average better than the baselines, DyHPO and DPL, except for the TaskSet benchmark family where DPL starts the best but ifBO improves with more budget.
Appendix F Per-task HPO Plots
In Section 3.1, we presented HPO results on each of these three benchmarks in a comprehensive form, averaging rank and normalized regrets across every task in the suite. These averages may hide / be susceptible to outliers. For completeness, Figures 12-17 provide regret plots for every task in the benchmark, averaged across the 10 seeds. We find that our method consistently performs on par, or better than the best previous best HPO method, especially in later stages of the search, without notable outliers.